OpenStudy (anonymous):

I have more awful maths hw D: What is the sum of the finite arithmetic series? (–15) + 0 + 15 + 30 + … + 195

OpenStudy (anonymous):

need a calculator essentially

OpenStudy (anonymous):

I tried using Wolfram Alpha and it says "(unable to determine general term)"

OpenStudy (anonymous):

what is the "..." stand for?

OpenStudy (jdoe0001):

hmmm what do you think is the common ratio?

OpenStudy (jdoe0001):

"..." ^ ellipses =P

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

every term after the last one is the last term + 15

OpenStudy (jdoe0001):

yes so what do you think is the common ratio?

OpenStudy (anonymous):

(–15) + 0 + 15 + 30 + 45 + 60 + 75 + 90 + 105 + 120 + 135 + 150 + 165 + 180 + 195 = 1350

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

can work with (1+2+3+...n) ) = (n)(n+1)/2 and a multiplier in front.

OpenStudy (anonymous):

a=-15,d=0-(-15)=15 l=a+(n-1)d find n,then \[Sn=\frac{ n }{2 }\left( a+l \right)\] or \[Sn=\frac{ n }{2 }\left\{ 2a+\left( n-1 \right)d \right\}\] Here l=195 calculate Sn.

OpenStudy (anonymous):

what??? i thought it was + 15?

OpenStudy (jdoe0001):

sum of an arithmetic sequence => \(\Large \bf S_{\color{red}{ n}}=\cfrac{{\color{red}{ n}}}{2}(a_1+a_{\color{red}{ n}}) \)

OpenStudy (anonymous):

Yes, 1350

OpenStudy (anonymous):

um, i don't know how to do that exactly, but i can try

OpenStudy (anonymous):

oh, it is right?

OpenStudy (anonymous):

yeah, it is right thanks mate :D wish i could give you more than just a medal